OpenStudy (anonymous):

Find what value of x makes this equation true... 0.6x+0.8=1.4

OpenStudy (blurbendy):

start by subtracting 0.8 from both sides

OpenStudy (blurbendy):

or realize that SOMETHING times 0.6 + 0.8 = 1.4

OpenStudy (anonymous):

What both side?

OpenStudy (blurbendy):

subtract 0.8 from both sides of the equation

OpenStudy (anonymous):

Okay:)

OpenStudy (blurbendy):

what did you get?

OpenStudy (anonymous):

Okay so I got -0.2 and 0.6

OpenStudy (blurbendy):

We start with: 0.6x + 0.8 = 1.4 then subtract 0.8 from both sides to get 0.6x = 0.6 because 1.4 - 0.8 = 0.6 understand?

OpenStudy (anonymous):

Oh okay so the value of x is 0.6

OpenStudy (blurbendy):

no, then we have to divide both sides by 0.6 x = 1 because 0.6 / 0.6 = 1

OpenStudy (anonymous):

O.m.g I totally forgot that!! Hah okay so it's one okay is that all??

OpenStudy (blurbendy):

yup, that's it

OpenStudy (anonymous):

The value of X that Will make the equation true is when x=1

OpenStudy (anonymous):

Oh okay thanks for your help!

OpenStudy (blurbendy):

no problem, anytime!

OpenStudy (toxicsugar22):

1